solve by substitution x+2y=8 and x+3y=12
Answer by jim_thompson5910(35256)

Hint: Solve for x in equation 1 to get x = 8-2y

Now plug in x = 8-2y into x+3y=12 to get 8-2y + 3y = 12


From here, solve for y. Once you have a value for y, use it to find x.